OpenStudy (anonymous):

One number is twelve more than twice another. If their sum is decreased by nine, the result is fifteen. Find the numbers. (Enter solutions from smallest to largest.)

OpenStudy (anonymous):

Let x = 1st number Let y = 2nd number x = 2y + 12 (x + y)-9 = 15 Which is... x = 2y + 12... 1 x + y = 24... 2 Plugging 1 into 2... 2y + 12 + y = 24 Therefore, y is 4 And, x is 20

OpenStudy (anonymous):

thank u